1.7t+8-1.62t=0.4t-.32+8
Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

1.7t+8-1.62t=0.4t-.32+8
0.08t + 8 = 0.4t + 7.68
-0.32t = - 0.32
t = 1
Hope this helps
:-)

Answer by stanbon(75887)

1.7t+8-1.62t = 0.4t-.32+8
---------------
Subtract 8 from both sides::
-------------
1.7t -1.62t - 0.4t = -0.32
------
-0.32t = -0.32
------
t = 1
===========
Cheers,
Stan H.
===============